# AFABench: A Generic Framework for Benchmarking Active Feature Acquisition

### v1

@@ Abstract.
In many real-world scenarios, acquiring all features of a data instance can be expensive or impractical due to monetary cost, latency, or privacy concerns. Active Feature Acquisition (AFA) addresses this challenge by dynamically selecting a subset of informative features for each data instance, trading predictive performance against acquisition cost. While numerous methods have been proposed for AFA, ranging from greedy information-theoretic strategies to non-myopic reinforcement learning approaches, fair and systematic evaluation of these methods has been hindered by the lack of standardized benchmarks. In this paper, we introduce AFABench, the first benchmark framework for AFA. Our benchmark includes a diverse set of synthetic and real-world datasets, supports a wide range of acquisition policies, and provides a modular design that enables easy integration of new methods and tasks. We implement and evaluate representative algorithms from all major categories, including static, greedy, and reinforcement learning-based approaches. To test the lookahead capabilities of AFA policies, we introduce a novel synthetic dataset, AFAContext, designed to expose the limitations of greedy selection. Our results highlight key trade-offs between different AFA strategies and provide actionable insights for future research. The benchmark code is available at: https://github.com/Linusaronsson/AFA-Benchmark.
@@ 1. Introduction
In many real-world applications, acquiring feature values of data instances can be costly. In healthcare, different medical tests can have monetary cost, latency, or privacy concerns. In recommender systems, querying user preferences can impose a cognitive burden or intrude on privacy. In such cases, acquiring the full set of features for every data instance may be infeasible or inefficient. Active Feature Acquisition (AFA) addresses this challenge by learning to dynamically select, for each data instance, a small subset of informative features to acquire, optimizing predictive performance under a constrained acquisition budget.
@@
Despite its relevance and importance, AFA has received significantly less attention than adjacent topics such as active learning (Settles, 2009). In active learning, all features are typically available, and the goal is to selectively acquire costly labels. In contrast, AFA assumes that both labels and features are available during training, but only partial features can be acquired at test time. AFA is also different from static feature selection methods, which fix a global subset of features regardless of the test instance. For comprehensive reviews of traditional feature selection methods, we refer the reader to (Guyon and Elisseeff, 2003; Li et al., 2018; Cai et al., 2018).
A recent survey by Rahbar et al. (2025) categorizes AFA methods into several paradigms. In general, the literature can be grouped into (i) greedy methods, which acquire features one by one based on expected utility (e.g., conditional mutual information), and (ii) non-greedy methods, typically framed via reinforcement learning, which learn acquisition policies that maximize long-term rewards rather than immediate gain.
Among greedy approaches, a common theme is to select features that maximize the expected information gain. Generative methods such as EDDI (Ma et al., 2019) estimate this using partial variational autoencoders that model the joint distribution over features and targets. In contrast, discriminative methods directly estimate expected improvements in prediction loss (e.g., (Covert et al., 2023; Gadgil et al., 2024)), avoiding the need to model the entire feature distribution. Non-greedy methods model the AFA problem as a Markov Decision Process (MDP), allowing feature acquisition to be treated as sequential decision-making under budget constraints. Model-free reinforcement learning methods such as JAFA (Shim et al., 2018) directly optimize acquisition policies, while model-based methods such as ODIN (Zannone et al., 2019) leverage generative models to simulate feature acquisition trajectories. These methods can learn lookahead strategies that outperform greedy selection in certain regimes.
@@
Although this body of work has grown considerably, progress is hindered by the lack of a standardized and unified evaluation framework. Most prior methods are evaluated in isolation using inconsistent datasets, model architectures, or acquisition costs. To our knowledge, no prior work has provided a dedicated benchmark for the AFA setting. This makes it difficult to perform fair comparisons or gain generalizable insights.
@@
This paper introduces the first benchmark for Active Feature Acquisition. Our framework provides a unified and extensible platform for comparing AFA methods under controlled and fair conditions. It supports a diverse set of synthetic and real-world datasets and includes representative algorithms from all major categories: greedy (generative and discriminative), non-greedy (model-free and model-based RL), and static baselines. More importantly, our framework is designed to be modular and easy to extend, allowing researchers to add new methods, acquisition strategies, and datasets with minimal engineering overhead. This makes our framework a valuable tool for both researchers developing new AFA techniques and practitioners seeking to apply AFA in domain-specific tasks.
@@
  • We implement and evaluate representative methods from all major paradigms, including greedy (generative and discriminative), reinforcement learning-based (model-free and model-based), and static feature selection.
  • We propose a novel synthetic dataset, AFAContext, that highlights the limitations of greedy strategies and provides a testbed for evaluating non-myopic acquisition policies.
@@ 2.1. Notation
Let $\mathbf{x}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{d}\}\in\mathbb{R}^{d}$ denote a data instance with $d$ features, and let $\mathbf{y}$ be the corresponding response variable in a supervised learning setting. We assume that $(\mathbf{x},\mathbf{y})$ is distributed according to a joint data distribution $p(\mathbf{x},\mathbf{y})$. Throughout, we use bold symbols (e.g., $\mathbf{x}$, $\mathbf{y}$) to represent random variables, and their non-bold counterparts (e.g., $x$, $y$) to denote specific realizations sampled from the distribution, i.e., $x,y\sim p(\mathbf{x},\mathbf{y})$. For any subset $S\subseteq[d]\triangleq\{1,\ldots,d\}$, we write $\mathbf{x}_{S}=\{\mathbf{x}_{i}\mid i\in S\}$ to denote the corresponding subset of features.
@@ 2.2. Active Feature Acquisition
As discussed in Section 1, the AFA problem appears in many variants. In this section, we describe a common formulation adopted for our benchmark.
In many machine learning settings, acquiring feature $i\in[d]$ incurs a monetary, latency, or privacy cost. AFA addresses this challenge by jointly (i) maximizing predictive performance and (ii) minimizing the total cost of queried features. Unlike static feature selection, which fixes the same feature subset for every data instance, AFA allows the set of acquired features to vary from one data instance to another, yielding a more flexible and realistic approach.
We therefore learn two functions: (i) a predictor $f(x_{S})$ that can make a prediction using any observed subset $S\subseteq[d]$; and (ii) an acquisition policy $\pi(x_{S})\in[d]\setminus S$, i.e., a function that, given the currently observed features $x_{S}$, returns the index of the next feature to reveal. They are optimized jointly via
(1) $\min_{f,\pi}\;\mathbb{E}_{p(\mathbf{x},\mathbf{y})}\!\left[\ell\!\bigl{(}f(\{\mathbf{x}_{a_{t}}\}_{t=1}^{b}),\,\mathbf{y}\bigr{)}\right],$
where $a_{n}=\pi(\{\mathbf{x}_{a_{t}}\}_{t=1}^{n-1})$ for $n=1,\dots,b$ and $b$ is a hard budget limiting how many feature values we may acquire per example. Eq. (1) adopts two simplifying assumptions: (i) all features incur the same acquisition cost, and (ii) every data instance receives an identical budget $b$. Consequently, exactly $b$ features are selected for each example. Although prior work has relaxed one or both of these assumptions, we retain them to ensure a fair and unbiased benchmark. A detailed rationale for this choice is provided in Section 3.4. However, it is straightforward to extend our framework to include these cases.

### v3

@@ Abstract.
In many real-world scenarios, acquiring all features of a data instance can be expensive or impractical due to monetary cost, latency, or privacy concerns. Active Feature Acquisition (AFA) addresses this challenge by dynamically selecting a subset of informative features for each data instance, trading predictive performance against acquisition cost. While numerous methods have been proposed for AFA, ranging from myopic information-theoretic strategies to non-myopic reinforcement learning approaches, fair and systematic evaluation of these methods has been hindered by a lack of standardized benchmarks. In this paper, we introduce AFABench, the first benchmark framework for AFA. Our benchmark includes a diverse set of synthetic and real-world datasets, supports a wide range of acquisition policies, and provides a modular design that enables easy integration of new methods and tasks. We implement and evaluate representative algorithms from all major categories, including static, myopic, and reinforcement learning-based approaches. To test the lookahead capabilities of AFA policies, we introduce a novel synthetic dataset, CUBE-NM, designed to expose the limitations of myopic selection. Our results highlight key trade-offs between different AFA strategies and provide actionable insights for future research. The benchmark code is available at: https://github.com/Linusaronsson/AFA-Benchmark.
@@ 1. Introduction
In many real-world applications, acquiring feature values of data instances can be costly. In healthcare, different medical tests can have monetary cost, latency, or privacy concerns (Turney, 1995). In recommender systems, querying user preferences can impose a cognitive burden or intrude on privacy (Jeckmans et al., 2013). Similar constraints arise in sensor systems and robotics, where measurements require power, time, or motion (Krause and Guestrin, 2009; Lauri et al., 2023). In such cases, acquiring the full set of features for every data instance may be infeasible or inefficient. Active Feature Acquisition (AFA) addresses this challenge by learning to dynamically select, for each data instance, a small subset of informative features to acquire, optimizing predictive performance under a constrained acquisition budget.
@@
Despite its relevance and importance, AFA has until recently received less attention than adjacent topics such as active learning (Settles, 2009). In active learning, all features are typically available, and the goal is to selectively acquire costly labels. In contrast, AFA assumes that both labels and features are available during training, but only partial features can be acquired at test time. AFA is also different from static feature selection methods, which fix a global subset of features regardless of the test instance. For comprehensive reviews of traditional feature selection methods, we refer the reader to (Guyon and Elisseeff, 2003; Li et al., 2018; Cai et al., 2018).
A recent survey by Aronsson et al. (2026) categorizes AFA methods into several paradigms. At a high level, the literature can be divided into (i) myopic methods, which select features by maximizing a one-step utility without explicit long-term planning, and (ii) non-myopic methods, which optimize long-term utility and are often formulated within a reinforcement learning (RL) framework. By learning policies that account for future acquisitions, non-myopic methods can yield lookahead strategies that outperform myopic selection in certain regimes. Section˜4.2 details the AFA methods considered in this work. A central question in this benchmark is whether the additional computational cost of non-myopic planning is justified relative to simpler myopic heuristics.
@@
Although this body of work has grown considerably in recent years, progress is hindered by the lack of a standardized and unified evaluation framework. Most existing methods are evaluated in isolation using inconsistent datasets, model architectures, or acquisition costs. To our knowledge, no prior work has provided a dedicated benchmark for the AFA setting. This makes it difficult to perform fair comparisons or gain generalizable insights.
@@
This paper introduces the first benchmark for AFA. We present a unified and extensible framework for comparing AFA methods under controlled, fair, and reproducible conditions. The benchmark supports a diverse collection of synthetic and real-world datasets and includes representative algorithms across the main methodological categories: myopic approaches (both generative and discriminative), non-myopic approaches (including model-free RL, model-based RL, and non-RL variants), as well as static baselines. Crucially, the framework is modular and easy to extend, enabling researchers to incorporate new methods, acquisition strategies, and datasets with minimal engineering effort. This makes our framework a valuable tool for both researchers developing new AFA techniques and practitioners seeking to apply AFA in domain-specific tasks.
@@
  • We implement and evaluate representative methods from all major paradigms, including myopic methods (generative and discriminative), non-myopic methods based on RL (model-free and model-based), non-myopic methods without explicit RL, and static feature selection.
  • We propose a novel synthetic dataset, Cube-NM, that highlights the limitations of myopic strategies and provides a testbed for evaluating non-myopic acquisition policies (which we justify theoretically, see Theorem˜4.1).
@@ 2.1. Notation
Let $\mathbf{x}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{d}\}\in\mathcal{X}$ denote a data instance with $d$ features, and let $\mathbf{y}\in\mathcal{Y}$ be the corresponding response variable in a supervised learning setting. We assume that $(\mathbf{x},\mathbf{y})$ is distributed according to a joint data distribution $p(\mathbf{x},\mathbf{y})$. Throughout, we use bold symbols (e.g., $\mathbf{x}$, $\mathbf{y}$) to represent random variables, and their non-bold counterparts (e.g., $x$, $y$) to denote specific realizations sampled from the distribution, i.e., $x,y\sim p(\mathbf{x},\mathbf{y})$. In practice, each $\mathbf{x}_{i}$ may correspond to a single feature (as is typical for tabular data), but it may also represent a group of features, for example a patch of pixels in an image. See Section 3 for details. For any subset $S\subseteq[d]\triangleq\{1,\ldots,d\}$, we write $\mathbf{x}_{S}=\{\mathbf{x}_{i}\mid i\in S\}\in\mathcal{X}_{S}$ to denote the corresponding subset of features. For an instance $x$, acquiring feature $i\in[d]$ reveals $x_{i}$ and incurs the cost $c_{i}\in\mathbb{R}^{+}$. In addition, we denote $c(S)\triangleq\sum_{i\in S}c_{i}$ for $S\subseteq[d]$.
@@ 2.2. Active Feature Acquisition
As discussed in Section 1, the AFA problem appears in many variants. In this section, we describe a common formulation adopted for our benchmark (Aronsson et al., 2026). In AFA, one learns (possibly jointly) two functions: (i) a predictor $f(S,x_{S})$ that produces an output given any observed subset $S\subseteq[d]$ (typically implemented via feature masking, see, e.g., (Aronsson et al., 2026)), and (ii) an acquisition policy $\pi(S,x_{S})\in\mathcal{A}_{B}(S)$, where
(1) $\mathcal{A}_{B}(S)\triangleq\{\,a\in U:c(S)+c_{a}\leq B\,\}\cup\{\texttt{STOP}\},$
is the set of allowed actions after acquiring $S$. Here, $U=[d]\setminus S$ denotes the set of unobserved features, and $B$ is a per-instance budget on cumulative feature cost. In practice, the policy often outputs a distribution over actions. If $\pi$ ouputs a feature acquisition action $a\in U$, we incur cost $c_{a}$ and observe $x_{a}$. We then update $S$ and $U$ as follows: $S\leftarrow S\cup\{a\}$ and $U\leftarrow U\setminus\{a\}$. If instead $\pi$ selects STOP, we output $\hat{y}=f(S,x_{S})$ and terminate. Let $\pi[x]\subseteq[d]$ denote the final set of acquired feature indices when the policy stops on instance $x$. Given this, the optimization objective for AFA is shown below.
(2) $\min_{f,\pi}\mathbb{E}_{\mathbf{x},\mathbf{y}}\mathbb{E}_{\pi}\left[\ell\big(f(\mathbf{x}_{\pi[\mathbf{x}]}),\mathbf{y}\big)+\alpha c(\pi[\mathbf{x}])\right]$
$\quad\text{s.t.}\quad c(\pi[x])\leq B\;\;\text{for all }x\in\mathcal{X}.$
where $\alpha\geq 0$ controls the cost–accuracy trade-off and $\ell$ is a loss function. The term $\alpha\,c(\pi[\mathbf{x}])$ allows the acquisition cost to vary across instances, enabling the policy to spend more resources on instances that are harder to predict.
Hard vs. soft budget. From the above, $\alpha$ induces a soft budget and is therefore referred to as a soft-budget parameter. By contrast, $B$ imposes a hard budget. A soft budget is beneficial because it adapts the cost spent per instance. However, selecting an appropriate value of $\alpha$ can be unintuitive in many applications (as discussed by, e.g., (Janisch et al., 2020)). A hard budget $B$ is more natural when resources are strictly limited per instance and is often easier to interpret. Its drawback is that some instances may require acquiring features whose total cost exceeds $B$ to achieve accurate predictions.
A note on terminology. Throughout the paper, the term AFA method refers to the predictor $f$ and the policy $\pi$ induced by the method. Some methods learn $f$ and $\pi$ jointly, for instance by training $f$ on the states $(S,x_{S})$ encountered under $\pi$ during learning. Other methods instead pretrain $f$ on randomly sampled states $(S,x_{S})$ and then use this predictor when training $\pi$ (optionally continuing to update $f$ using states encountered under $\pi$). This distinction is discussed further in Section˜3.2.
2.3. Scope of the Benchmark
Most prior work on AFA considers two special cases of (LABEL:eq:afaopt): (i) the hard-budget setting, where $\alpha=0$ (only a hard budget), and (ii) the soft-budget setting, where $B=c([d])$ (all features are, in principle, acquirable for every instance, so only the soft-budget term matters). Following prior work, we also focus on these special cases in our benchmark.